The range of the function f(x)=x+5 is {7,9}. What is the function's domain?
Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39617)

Solve for x and substitute the given range elements for f(x).

Answer by ikleyn(52781)

.
One bound of the function domain is 

x1 + 5 = 7  --->  x1 = 7-5 = 2.

Other bound of the function domain is 

x2 + 5 = 9  --->  x2 = 9-5 = 4.

The domain is [2,4].

When you write the info for the range, do not use these parentheses "{" and "}".
They are from other story.

Use these parentheses "(" and ")" or "[" and "]", or their combinations.